OpenStudy (anonymous):

can you help me.., why The scattered peak of Davisson and Germer 's electron scattering occurs for 50o ?

OpenStudy (anonymous):

it is just because of the wave nature of electron as the braggs law give 1/lamda=n/2d sinx=p/h=sqt[2meV]/h and as the d=const, h=const so sinx is directly prop to the square root of V so at 54V angle comes out to be 50.simple:)

OpenStudy (anonymous):

So what the fixed value of the collector current in the detector occurs for 50o? and for 70o?

OpenStudy (anonymous):

can you tell me @Aperogalics ?

OpenStudy (anonymous):

wait plz.........u r in the que:)

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

To measure the number of electrons that were scattered at different angles, an electron detector that could be moved on an arc path about the crystal was used. The detector was designed to accept only elastically scattered electrons it is just a sensitive detector which works like galvanometer it just depends how much intensity of electrons is tere....:)

OpenStudy (anonymous):

how the mathematical calculations? ?

OpenStudy (anonymous):